In a recent test of its braking system, a Volkswagen Passat traveling at 31.2 m/s came to a full stop after an average negative acceleration of magnitude 2.10 m/s2.

(a) How many revolutions did each tire make before the car comes to a stop, assuming the car did not skid and the tires had radii of 0.350 m?

(b) What was the angular speed of the wheels (in rad/s) when the car had traveled half the total stopping distance?


1
Expert's answer
2021-10-25T10:04:32-0400

a) According to the kinematic equation 4 (see https://www.khanacademy.org/science/physics/one-dimensional-motion/kinematic-formulas/a/what-are-the-kinematic-formulas), the distance travelled by the car to full stop is (taking the final speed, "v", to be 0):


"d =\dfrac{v_0^2}{2a}"

where "v_0 = 31.2m\/s" is the initial speed, "a = 2.10m\/s^2" is its acceleration.

In one turn tire travels the distance equal to the length of its circumference:


"l = 2\pi r"

where "r = 0.350m" is its radius. Thus, the total number of turns is:


"N = \dfrac{d}{l} = \dfrac{ v_0^2}{4\pi ra}\\
N = \dfrac{(31.2m/s)^2}{4\pi \cdot 0.35m\cdot2.1m/s^2} \approx 105"

b) The angular speed is given as follows:


"\omega = \dfrac{v}{r}"

where "v" is the linear speed. Using the same kinematic equation, find the speed at "d\/2":


"v = \sqrt{v_0^2 - ad} = \sqrt{v_0^2 - a\cdot \dfrac{v_0^2}{2a}} = \dfrac{v_0}{\sqrt{2}}"

where '-' sign was used, because the acceleration is negative. Thus, obtain:


"\omega = \dfrac{v_0}{r\sqrt{2}} = \dfrac{31.2m/s}{0.35m\cdot \sqrt{2}} \approx 63.0rad/s"

Answer. a) 105, b) 63.0 rad/s.